OpenStudy (anonymous):

Can someone please explain to me how to do this? 6^4 Square root 9 + 2^3 square root 9 + 2^4 square root 9 - 3^3 square root 9

OpenStudy (anonymous):

i love ur name!

mathslover (mathslover):

Hello @Musical.Panda

mathslover (mathslover):

Let me make the question look nice . \[\large{6^4 \sqrt{9} + 2^3 \sqrt{9} + 2^4 \sqrt{9} - 3^3 \sqrt{9}}\] is this what we have ?

OpenStudy (anonymous):

Yea, that's it.

mathslover (mathslover):

So what is \(\sqrt{9}\) ?

OpenStudy (anonymous):

3

mathslover (mathslover):

great so we have : \[\large{6^4 (3) + 2^3 (3) + 2^4 (3) - 3^3 (3)}\] \[\large{3 ( 6^4 + 2^3 + 2^4 - 3^3 ) }\]

mathslover (mathslover):

I think you will have to use some calculator and find the values and then you will arrive with your final answerr

OpenStudy (anonymous):

ok thanks :)

mathslover (mathslover):

Also welcome to openstudy

OpenStudy (anonymous):

Thanks :)